# Better Estimation of the KL Divergence Between Language Models

### v2

@@ 5.1 Understanding the KL Estimators
$\delta_{\textsc{mc}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{ht}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{cv}}$ $6.28\pm 2.54$ $6.28\pm 1.13$ $6.28\pm 0.79$
$\delta_{\textsc{cv}1}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\delta_{\textsc{rb}}$ $6.76\pm\mathbf{0.11}$ $6.76\pm\mathbf{0.05}$ $6.76\pm\mathbf{0.03}$
Table 1: Estimated value $\pm$ empirical standard deviation of different estimators. When aggregating over prompts, $\delta_{\textsc{ht}}$ and $\delta_{\textsc{cv}}$ fail to significantly reduce the variance of $\delta_{\textsc{mc}}$. Rao–Blackwellized estimator, however, achieves the lowest standard deviation.
@@
In Tab. 1, we report the expected KL estimate along with the empirical standard deviation of different estimators evaluated at sample sizes $M=1,5,10$. To obtain these estimates, we compute each estimator using $M$ samples, repeating the process $4000/M$ times to estimate both the expected value and the standard deviation of the estimates. Our findings confirm that all estimators except one ($\delta_{\textsc{cv}},\alpha=1$), are unbiased and report an expected KL divergence of $6.76$. We also observe that the CV estimators fail to significantly reduce the variance of the standard Monte Carlo estimator. Importantly, Rao–Blackwellized Monte Carlo estimator achieves the lowest standard deviation and offers a more robust estimate compared to the MC estimator. Interestingly, when setting $\alpha=1$, the control variate Monte Carlo estimator exhibits both noticeable bias and high variance. To investigate this further, we next examine KL divergence estimates at the level of individual prompts.
Figure 2: Comparison of the Monte Carlo (MC) and Rao–Blackwellized (RB) estimators in the RLHF fine-tuning loop. We perform RLHF with each estimator $5$ times and plot the mean and standard deviation (in shades) of the average reward values and the KL at each fine-tuning step. We observe that the MC estimator is not as stable as the RB estimator and its performance varies significantly across different runs. However, RB estimator reliably offers a good balance between achieving low KL and high reward values in all runs.
In Fig. 1, we visualize the KL estimates for three different prompts: (left) a positive prompt, (middle) a neutral prompt, and (right) a negative adversarial prompt. The traces represent the average estimates from all the repetitions, while the shaded regions indicate the standard deviation. Except $\delta_{\textsc{cv}},\alpha=1$, all other estimators are unbiased and consistent. We observe that the $\delta_{\textsc{cv}}$ estimator exhibits a noticeable bias and high standard deviation when $\alpha=1$, i.e., when it is not set to its optimal value. The bias arises from numerical instability during the computation of $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1$. The high variance is due to large values of $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$. Specifically, for certain prompts, $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$ can be unbounded.
Figure 3: Standard deviation of different KL estimators across various prompts in the IMDB evaluation set, displayed on a logarithmic scale. The $\delta_{\textsc{cv}}$ estimator with $\alpha=1$ shows considerable instability for certain prompts, whereas the $\delta_{\textsc{rb}}$ estimator consistently yields the lowest standard deviation, indicating more stable estimates.
@@
Since the robustness of the estimators depends on the choice of the prompt, we further analyze their estimated standard deviations across all prompts. Fig. 3 presents a box plot of standard deviations (log scale) for each estimator. The $\delta_{\textsc{cv}}$ estimator with $\alpha=1$ shows significant instability for certain prompts, with numerous outliers indicating high variance. In contrast, the $\delta_{\textsc{mc}}$ and the standard $\delta_{\textsc{cv}}$ estimators exhibit comparable standard deviations. Notably, the Rao–Blackwellized estimator consistently achieves the lowest standard deviation, suggesting it provides the most stable estimates.
@@ 5.2 KL Estimation and RLHF Training Dynamics
A key application of KL estimation is in the RLHF training loop. From the previous experiment, we observed that the Rao–Blackwellized estimator significantly reduces the standard deviation of the Monte Carlo estimator. Therefore, it is natural to ask how this affects RLHF performance when this estimator is used in the training loop. The RLHF objective consists of two terms: (i) the expected rewards for samples generated by the language model $p_{\boldsymbol{\theta}}$, which in this case is the samples’ score under a sentiment classifier [^11] Specifically, we look at the logits of the positive class., and (ii) the negative KL divergence between the fine-tuned model $p_{\boldsymbol{\theta}}$ and the reference model $q$, which represents the language model before fine-tuning.
Figure 4: Compared to MC-trained models, RB-trained models appear on the Pareto front $76\%$ of the time. We plot the reward vs. KL Pareto front for each estimator to illustrate how RB more effectively constrains KL while maintaining competitive rewards.
@@
We compare the Monte Carlo and Rao–Blackwellized estimators for computing the gradient of the KL divergence term in the RLHF training loop. We use the RL algorithm [^12] We discuss a common mistake when Rao–Blackwellizing the KL estimator in trust-region algorithms in § F.2. proposed by Ahmadian et al. (2024), [^13] Specifically, we use the available implementation of this algorithm in the trl library (von Werra et al., 2020). which is an improved version of the reinforce algorithm (Williams, 1992). [^14] App. G reports the hyperparameters used for the algorithm. We track two metrics throughout fine-tuning: (i) the average reward associated with samples drawn from $p_{\boldsymbol{\theta}}$, and (ii) the KL divergence between $p_{\boldsymbol{\theta}}$ and $q$. The results are visualized in Fig. 2. The purple trace represents the training run where the $\delta_{\textsc{rb}}$ is used in the optimization loop to estimate the gradient of the KL divergence, while the blue trace represents the run using $\delta_{\textsc{mc}}$. The x-axis denotes the fine-tuning step, with the left plot showing the evolution of the average reward and the right plot displaying the KL divergence between $p_{\boldsymbol{\theta}}$ and $q$ over the course of fine-tuning. We repeat each experiment $5$ times and report the mean and standard deviation of each metric. Notably, the KL values in the right plot are estimated using the Rao–Blackwellized estimator. However, we observe the same overall trend when using the Monte Carlo estimator for evaluation.
@@
As illustrated in Fig. 2, the performance of the models trained using the MC estimator varies significantly across the $5$ experiments, resulting in a large standard deviation in both average rewards and the KL divergence. However, Rao–Blackwellized estimator consistantly achieves high rewards and reasonable KL values across all runs. This observation suggests that reducing the variance of the KL estimator makes the RLHF more reliable.
@@
Finally, we vary the KL coefficient, $\beta$, in $[0.01,0.1]$ range and fine-tune $18$ models with each estimator. For each estimator, we plot the Pareto frontier of average rewards versus KL divergence in Fig. 4, displaying models that do not appear on the Pareto front with reduced opacity. Overall, we find that fine-tuning with the RB estimator is more effective at achieving high rewards while maintaining a low KL divergence from the reference model. To quantify this effect, we compute the fraction of RB fine-tuned models that appear on the overall Pareto front—i.e., the frontier obtained when considering all models fine-tuned with either estimator. We then conduct a permutation test and report the results in Fig. 4. We find that $76\%$ of the points on the overall Pareto front come from RB fine-tuned models. Restricting to models with KL values below $5$, this fraction rises to $95\%$, with both results being statistically significant.
@@ 6 Conclusion
In this paper, we study the problem of estimating the KL divergence between language models. We provide a comprehensive formal analysis of various KL estimators, with a focus on their bias and variance. We introduce a Rao–Blackwellized Monte Carlo estimator, which is provably unbiased and has variance at most equal to that of the standard Monte Carlo estimator. This estimator applies the well-known Rao–Blackwellization technique to reduce the variance of the standard Monte Carlo method. Our empirical results show that the Rao–Blackwellized estimator significantly reduces the variance compared to the Monte Carlo estimator, while other estimators fail to achieve meaningful variance reduction or, in some cases, suffer from unbounded variance. Additionally, we find that using our proposed Rao–Blackwellized estimator makes RLHF more stable and produces models that more frequently lie on the Pareto frontier of reward versus KL, compared to models fine-tuned with the Monte Carlo estimator.
@@ Appendix A Horvitz–Thompson Estimation
When estimating the KL divergence between two language models $p$ and $q$, we have access not only to samples from $p$ but also to the probability assigned by $p$ to any string $\boldsymbol{y}$. This enables the use of a more informed estimator, which leverages these probabilities during its construction. Notably, this estimator is a specific instance of the Horvitz–Thompson (HT) estimator (Horvitz and Thompson, 1952), defined as []
$\delta_{\textsc{ht}}=\sum_{\boldsymbol{y}\in S}\frac{p(\boldsymbol{y})}{\pi_{S}(\boldsymbol{y})}\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}=\sum_{\boldsymbol{y}\in S}\frac{p(\boldsymbol{y})}{\pi_{S}(\boldsymbol{y})}f(\boldsymbol{y}),$ (30)
@@
$\delta_{\textsc{ht}}$ is an unbiased estimator of the KL divergence, i.e., $\operatorname*{\mathbb{E}}\big{[}\delta_{\textsc{ht}}\big{]}-\textsc{kl}(p\mid\mid q)=0$.
@@
$\operatorname*{\mathbb{E}}_{S}\Big{[}\delta_{\textsc{ht}}\Big{]}-\textsc{kl}(p\mid\mid q)$ $=\operatorname*{\mathbb{E}}_{S}\Big{[}\sum_{\boldsymbol{y}\in S}\frac{p(\boldsymbol{y})}{\pi_{S}(\boldsymbol{y})}\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}\Big{]}-\textsc{kl}(p\mid\mid q)$ (definition of $\delta_{\textsc{ht}}$) (31a)
@@
[…]
@@
Similar to the MC estimator, the HT estimator does not necessarily return a non-negative estimate of the KL. In principle, however, we should prefer Eq. 30 to Eq. 6a because it exploits more information—namely, the knowledge of $p$. Whether the HT estimator yields lower variance than the MC estimator depends on the sampling design used to construct $S$. In our experiments in App. G, we used the sampling-with-replacement design, where $\pi_{S}(\boldsymbol{y})=1-\left(1-p(\boldsymbol{y})\right)^{M}$. Compared to the MC estimator, we observed no significant reduction in variance in our experiments.
Appendix B Proof of Prop. 1
@@
First, we remind the reader that the control variate Monte Carlo estimator was defined in Eq. 8 as See 4
Note that $\{\boldsymbol{Y}^{(m)}\}_{m=1}^{M}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$. Next, we compute the bias and variance of this estimator as follows:
$\operatorname*{\mathbb{E}}\left[\delta_{\textsc{cv}}\right]-\textsc{kl}=\frac{1}{M}\sum_{m=1}^{M}\operatorname*{\mathbb{E}}\left[f(\boldsymbol{Y}^{(m)})\right]+\alpha\underbrace{(\operatorname*{\mathbb{E}}\left[g(\boldsymbol{Y}^{(m)})\right]-G)}_{=0}-\textsc{kl}=\operatorname*{\mathbb{E}}\left[f(\boldsymbol{Y})\right]-\textsc{kl}=0$ (32a)
$\mathop{\mathrm{Var}}\left[{\normalcolor\delta_{\textsc{cv}}}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f+\alpha(g-G)}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]+\frac{\alpha^{2}}{M}\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]+\frac{2\alpha}{M}\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right].$ (32b)
Appendix C Proof of Prop. 2
See 2
@@
[…]
@@
Appendix D Connection to Bregman Divergence
Bregman divergence (Bregman, 1967) is a measure of the difference between a convex function and its tangent plane. Let $F\colon\Omega\rightarrow\mathbb{R}$ be a strictly convex function, where $\Omega$ is a convex set. Bregman divergence associated with $F$ is defined as
$D_{F}(p,q)=F(p)-F(q)-\langle\nabla_{\boldsymbol{\theta}}F(q),p-q\rangle.$ (34)
@@ Appendix D Connection to Bregman Divergence
Let $F(p)=-\mathrm{H}(p)$, defined as $\mathrm{H}(p)=\sum_{\boldsymbol{y}\in\Sigma^{*}}-p(\boldsymbol{y})\log p(\boldsymbol{y})$. Then, the Bregman divergence $D_{F}(p,q)=\textsc{kl}(p\mid\mid q)$.
@@
$D_{F}(p,q)$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log p(\boldsymbol{y})-q(\boldsymbol{y})\log q(\boldsymbol{y})-\big{(}\log q(\boldsymbol{y})+1\big{)}\big{(}p(\boldsymbol{y})-q(\boldsymbol{y})\big{)}$ (35a)
$=\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\log p(\boldsymbol{y})-q(\boldsymbol{y})\log q(\boldsymbol{y})+q(\boldsymbol{y})\log q(\boldsymbol{y})-p(\boldsymbol{y})\log q(\boldsymbol{y})-p(\boldsymbol{y})+q(\boldsymbol{y})$ (35b)
$=\sum_{\boldsymbol{y}\in\Sigma^{*}}p(\boldsymbol{y})\Big{(}\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}+\frac{q(\boldsymbol{y})}{p(\boldsymbol{y})}-1\Big{)}$ (35c)

### v3

@@ 5.1 Analyzing the KL Estimators
$\mu_{\textsc{mc}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\mu_{\textsc{ht}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\mu_{\textsc{cv}1}$ $6.28\pm 2.54$ $6.28\pm 1.13$ $6.28\pm 0.79$
$\mu_{\textsc{cv}}$ $6.76\pm 0.16$ $6.76\pm 0.07$ $6.76\pm 0.05$
$\mu_{\textsc{rb}}$ $6.76\pm\mathbf{0.11}$ $6.76\pm\mathbf{0.05}$ $6.76\pm\mathbf{0.03}$
To determine the preferred response in each pair, we employ a binary sentiment classifier, [^15] Specifically, we use https://huggingface.co/lvwerra/distilbert-imdb. selecting the response with the higher probability of positive sentiment. Upon successful fine-tuning, $p_{\boldsymbol{\theta}}$ should assign a higher probability mass to movie reviews with positive sentiment while maintaining a low KL divergence with $q$. We then evaluate this KL divergence using our estimators to assess their reliability in measuring distributional shifts induced by fine-tuning.
Figure 1: Standard deviation of KL estimators across various prompts in the IMDB datasest.
We evaluate on 512 examples from the IMDB dataset. For each review, we randomly select a prefix length between 2 and 8 tokens and use it as the prompt. we then generate 4000 samples from $p_{\boldsymbol{\theta}}$ for each prompt. Using these samples, we compute the MC, control variate (CV), and Rao–Blackwellized (RB) estimators and estimate their standard We also implement the Horvitz–Thompson (HT) estimator; see App. A for details. The CV estimator, $\mu_{\textsc{cv}}$, is computed twice: once using the optimal $\alpha$ estimated from 1000 samples, and once with $\alpha=1$ to match the setup in Schulman [33].
@@
In Tab. 1, we report the expected KL estimate along with the empirical standard deviation of different estimators evaluated at sample sizes $M=1,5,10$. To obtain these estimates, we compute each estimator using $M$ samples, repeating the process $4000/M$ times to estimate both the expected value and the standard deviation of the estimates. Our findings confirm that all estimators except one ($\mu_{\textsc{cv}},\alpha=1$), are unbiased and report an expected KL divergence of $6.76$. We also observe that the CV estimators fail to significantly reduce the variance of the standard MC estimator. Importantly, the RB estimator achieves the lowest standard deviation and offers a more robust estimate compared to the MC estimator. Interestingly, we observe that the $\mu_{\textsc{cv}}$ estimator exhibits a noticeable bias and high standard deviation when $\alpha=1$, i.e., when it is not set to its optimal value. The bias arises from numerical instability during the computation of $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}-1$. The high variance is due to large values of $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$. Specifically, for certain prompts, $\mathop{\mathrm{Var}}\left[{\normalcolor g(\boldsymbol{Y})}\right]$ can be unbounded. We visualize the estimates for 33 example prompts in App. E.
@@
Since the robustness of the estimators depends on the choice of the prompt, we further analyze their estimated standard deviations across all prompts. Fig. 1 presents a box plot of standard deviations (in log scale) for each estimator. The $\mu_{\textsc{cv}}$ estimator with $\alpha=1$ shows significant instability for certain prompts, with numerous outliers indicating high variance. In contrast, the $\mu_{\textsc{mc}}$ and the standard $\mu_{\textsc{cv}}$ estimators exhibit comparable standard deviations. In particular, the Rao-Blackwellized estimator consistently achieves the lowest standard deviation, suggesting that it provides the most stable estimates.
@@ 5.2 KL Estimation and RLHF Training Dynamics
A key application of KL estimation is in the RLHF training loop. From the previous experiment, we observed that the RB estimator significantly reduces the standard deviation of the MC estimator. Therefore, it is natural to ask how this affects RLHF performance when this estimator is used in the training loop. The RLHF objective consists of two terms: (i) the expected rewards for samples generated by the language model $p_{\boldsymbol{\theta}}$, which in this case is the samples’ score under a sentiment classifier [^16] Specifically, we look at the logits of the positive class., and (ii) the negative KL divergence between the fine-tuned model $p_{\boldsymbol{\theta}}$ and the reference model $q$, which represents the language model before fine-tuning.
Figure 2: Comparison of the Monte Carlo (MC) and Rao–Blackwellized (RB) estimators in the RLHF fine-tuning loop. We perform RLHF with each estimator 55 times and plot the mean and standard deviation (in shades) of the average reward values and the KL at each fine-tuning step. We observe that the MC estimator is not as stable as the RB estimator and its performance varies significantly across different runs. However, RB estimator reliably offers a good balance between achieving low KL and high reward values in all runs.
We compare the MC and RB estimators for computing the gradient of the KL divergence term in the RLHF training loop. We use the RL algorithm171717§ D.2 discusses a common mistake when Rao–Blackwellizing the KL estimator in trust-region algorithms. proposed by Ahmadian et al. [2], [^18] Specifically, we use the available implementation of this algorithm in the trl library [40]. which is an improved verfaiion of the reinforce algorithm [41]. [^19] App. F reports the hyperparameters used for the algorithm.
Figure 3: Compared to models trained with MC esimator, models trained with RB appear on the Pareto front $78\%$ of the time.
First, we empirically test Thm. 5 by measuring the variance of the gradient norm. We sample $40$ prompts and compute the gradient of the KL divergence with respect to the model parameters using both the MC and RB estimators. To estimate variance empirically, we repeat this process 55 times. Both estimators are evaluated on the same prompts and model initializations to ensure a fair comparison. We find that the variance of the gradient norm estimated with the MC estimator is $59.90$, whereas with the RB estimator it is $45.44$. This corresponds to a $24.6\%$ reduction in variance, providing direct empirical evidence consistent with our theoretical motivation.
@@
We then proceed with using both estimators in RL fine-tuning. We track two metrics throughout fine-tuning: (i) the average reward associated with samples drawn from $p_{\boldsymbol{\theta}}$, and (ii) the KL divergence between $p_{\boldsymbol{\theta}}$ and $q$. The results are visualized in Fig. 2. The purple trace represents the training run where the $\mu_{\textsc{rb}}$ is used in the optimization loop to estimate the gradient of the KL divergence, while the blue trace represents the run using $\mu_{\textsc{mc}}$. The x-axis denotes the fine-tuning step, with the left plot showing the evolution of the average reward and the right plot displaying the KL divergence between $p_{\boldsymbol{\theta}}$ and $q$ over the course of fine-tuning. We repeat each experiment 55 times and report the mean and standard deviation of each metric. Notably, the KL values in the right plot are estimated using the RB estimator. However, we observe the same overall trend when using the MC estimator for evaluation.
@@
As illustrated in Fig. 2, the performance of the models trained using the MC estimator varies significantly across the 55 experiments, resulting in a large standard deviation in both average rewards and the KL divergence. However, RB estimator consistantly achieves high rewards and reasonable KL values across all runs. This observation suggests that the RB estimator makes RLHF runs more stable.
@@
Finally, we vary the KL coefficient, $\beta$, in $[0.01,0.1]$ range and fine-tune $18$ models with each estimator. For each estimator, we plot the Pareto frontier of average rewards versus KL divergence in Fig. 3, displaying models that do not appear on the Pareto front with reduced opacity. Overall, we find that fine-tuning with the RB estimator is more effective at achieving high rewards while maintaining a low KL divergence from the reference model. To quantify this effect, we compute the fraction of RB fine-tuned models that appear on the overall Pareto front—i.e., the frontier obtained when considering all models fine-tuned with either estimator. We then conduct a permutation test and report the results in Fig. 3. We find that $78\%$ of the points on the overall Pareto front come from RB fine-tuned models. Restricting to models with KL values below 55, this fraction rises to $95\%$, with both results being statistically significant.
@@ 6 Conclusion
In this paper, we study the problem of estimating the KL divergence between language models. We provide a comprehensive formal analysis of various KL estimators, with a focus on their bias and variance. We introduce the RB estimator, which is provably unbiased and has variance at most equal to that of the standard NC estimator. This estimator applies the well-known Rao–Blackwellization technique to reduce the variance of the standard MC method. Our empirical results show that the RB estimator significantly reduces the variance compared to the MC estimator, while other estimators fail to achieve meaningful variance reduction or, in some cases, suffer from unbounded variance. Additionally, we find that using our proposed RB estimator makes RLHF more stable and produces models that more frequently lie on the Pareto frontier of reward versus KL, compared to models fine-tuned with the MC estimator.
@@
Limitations
In our RLHF experiments, evaluating the variance of our estimator and comparing it to existing methods requires training a large number of models. For instance, the significance test in § 5.2 involves training 36 models. Due to limited computational resources, we used the controlled-generation task as a testbed. Given the strength of both our theoretical and empirical results, we hope future work will adopt the Rao–Blackwellized estimator and apply it to larger language models and a wider variety of RL-based approaches to LM alignment.
@@ Appendix A Horvitz–Thompson Estimation
When estimating the KL divergence between language models $p$ and $q$, we have access not only to samples from $p$ but also to the probability assigned by $p$ to any string $\boldsymbol{y}$. This enables the use of a more informed estimator, which leverages these probabilities during its construction. Notably, this estimator is a specific instance of the Horvitz–Thompson (HT) estimator [16], defined as []
$\mu_{\textsc{ht}}=\sum_{\boldsymbol{y}\in S}\frac{p(\boldsymbol{y})}{\pi_{S}(\boldsymbol{y})}\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}=\sum_{\boldsymbol{y}\in S}\frac{p(\boldsymbol{y})}{\pi_{S}(\boldsymbol{y})}f(\boldsymbol{y}),$ (22)
@@
$\mu_{\textsc{ht}}$ is an unbiased estimator of the KL divergence, i.e., $\operatorname*{\mathbb{E}}\big[\mu_{\textsc{ht}}\big]=\textsc{kl}(p\mid\mid q)$.
@@
$\operatorname*{\mathbb{E}}_{S}\Big[\mu_{\textsc{ht}}\Big]-\textsc{kl}(p\mid\mid q)$
$=\operatorname*{\mathbb{E}}_{S}\Big[\sum_{\boldsymbol{y}\in S}\frac{p(\boldsymbol{y})}{\pi_{S}(\boldsymbol{y})}\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}\Big]-\textsc{kl}(p\mid\mid q)$ (definition of $\mu_{\textsc{ht}}$) (23a)
@@
$=\sum_{\boldsymbol{y}\in\Sigma^{*}}\frac{p(\boldsymbol{y})}{\cancel{\pi_{S}(\boldsymbol{y})}}\log\frac{p(\boldsymbol{y})}{q(\boldsymbol{y})}\,\cdot\,\cancel{\pi_{S}(\boldsymbol{y})}-\textsc{kl}(p\mid\mid q)$ (definition of $\pi_{S}(\boldsymbol{y})$) (23d)
@@
Similar to the MC estimator, the HT estimator does not necessarily return a non-negative estimate of the KL. In principle, however, we should prefer Eq. 22 to Eq. 4 because it exploits more information—namely, the knowledge of pp. Whether the HT estimator yields lower variance than the MC estimator depends on the sampling design used to construct $S$. In our experiments in App. E, we used the sampling-with-replacement design, where $\pi_{S}(\boldsymbol{y})=1-\left(1-p(\boldsymbol{y})\right)^{M}$. Compared to the MC estimator, we observed no significant reduction in variance in our experiments.
Appendix B Control Variate Monte Carlo Estimation
@@
Recall that the control variate Monte Carlo estimator is defined as See 7 Note that $\{\boldsymbol{Y}^{(m)}\}_{m=1}^{M}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny i.i.d.}}}{{\sim}}}p$. First, we look at the expected value of the estimator:
$\operatorname*{\mathbb{E}}\left[\mu_{\textsc{cv}}\right]=\frac{1}{M}\sum_{m=1}^{M}\operatorname*{\mathbb{E}}\left[f(\boldsymbol{Y}^{(m)})\right]+\alpha\cdot\underbrace{(\operatorname*{\mathbb{E}}\left[g(\boldsymbol{Y}^{(m)})\right]-G)}_{=0}=\operatorname*{\mathbb{E}}\left[f\right]$ (24a)
Therefore, it is unbiased. Next, we manipulate the variance
$\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{cv}}}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f+\alpha\cdot(g-G)}\right]=\frac{1}{M}\mathop{\mathrm{Var}}\left[{\normalcolor f}\right]+\frac{\alpha^{2}}{M}\mathop{\mathrm{Var}}\left[{\normalcolor g}\right]+\frac{2\alpha}{M}\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right],$ (24b)
where the second equality above stems from well-known variance identities. [^20] Let $X$ and $Y$ be real-valued random variables, and let $a$ be a real scalar. Then, the following hold: $\mathop{\mathrm{Var}}\left[{\normalcolor X+Y}\right]=\mathop{\mathrm{Var}}\left[{\normalcolor X}\right]+\mathop{\mathrm{Var}}\left[{\normalcolor Y}\right]+2\mathop{\mathrm{Cov}}\left[{\normalcolor X,Y}\right]$, $\mathop{\mathrm{Var}}\left[{\normalcolor aX}\right]=a^{2}\mathop{\mathrm{Var}}\left[{\normalcolor X}\right]$, and $\mathop{\mathrm{Var}}\left[{\normalcolor X+a}\right]=\mathop{\mathrm{Var}}\left[{\normalcolor X}\right]$.
Proposition 7.
Provided that the target KL divergence is finite, the estimator $\mu_{\textsc{cv}}$ with $\alpha=1$ is nonnegative with probability one. Furthermore, $\alpha=1$ is the only value for which we can guarantee nonnegativity of $\mu_{\textsc{cv}}$.
Proof.
To prove that $\mu_{\textsc{cv}}$ is nonnegative with probability one when $\alpha=1$, we will prove that each term $\log\frac{p(\boldsymbol{Y}^{(m)})}{q(\boldsymbol{Y}^{(m)})}+\alpha\cdot\left(\frac{q(\boldsymbol{Y}^{(m)})}{p(\boldsymbol{Y}^{(m)})}-1\right)$ in the summation that defines $\mu_{\textsc{cv}}$ is nonnegative.
Let $r=\frac{q(\boldsymbol{Y}^{(m)})}{p(\boldsymbol{Y}^{(m)})}$. We identify $\alpha$ for which $\log\frac{1}{r}+\alpha\cdot(r-1)\geq 0$ holds for all r>0r>0.212121Note that $r>0$ whenever this term is finite. Equivalently, we want to find $\alpha$ values for which:
$\inf_{r>0}-\log r+\alpha\cdot(r-1)\geq 0.$ (25a)
Next, we prove that $\alpha=1$ is the only value satisfies the inequality Eq. 25a. To see why, consider the first-order optimality conditions for the minimization over $r$. These conditions are necessary and sufficient as $\log r+\alpha\cdot(r-1)$ is convex for all $\alpha$ over the region $r>0$. Solving for $r$ such that the derivative is zero gives us
$0$ $=\frac{\partial}{\partial r}\left[-\log r+\alpha\cdot(r-1)\right]$ (25b)
$\iff 0$ $=-1/r+\alpha$ (25c)
$\iff r$ $=1/\alpha.$ (25d)
Plugging that value allows us to simplify away the minimization:
$0$ $\leq\inf_{r>0}-\log r+\alpha\cdot(r-1)$ (25e)
$=-\log(1/\alpha)+\alpha\cdot(1/\alpha-1)$ (25f)
$=\log(\alpha)+1-\alpha.$ (25g)
For each term in $\mu_{\textsc{cv}}$ to be nonnegative, we need $\log(\alpha)+1-\alpha\geq 0$. However, the function $\log(\alpha)+1-\alpha$ is strictly concave and nonpositive for all $\alpha>0$, attaining zero only at $\alpha=1$. Thus, when $\alpha=1$ every term in $\mu_{\textsc{cv}}$ is nonnegative. When $\alpha\neq 1$, some terms may be nonnegative, leading to the possibility that $\mu_{\textsc{cv}}$ may be negative.
Proposition 8.
Given the random variable $\boldsymbol{Y}\sim p$, let $f(\boldsymbol{Y})=\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}$ and $g(\boldsymbol{Y})=\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}$. We prove $\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]$ is zero if and only if $p=q$.
@@
$=\operatorname*{\mathbb{E}}\left[\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}-\cancel{\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}}-\operatorname*{\mathbb{E}}\left[\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}\right]\frac{q(\boldsymbol{Y})}{p(\boldsymbol{Y})}+\cancel{\operatorname*{\mathbb{E}}\left[\log\frac{p(\boldsymbol{Y})}{q(\boldsymbol{Y})}\right]}\right]$ (26b)
@@
Since each KL divergence is non-negative and equals zero if and only if the two distributions coincide almost everywhere, we have,
$\mathop{\mathrm{Cov}}\left[{\normalcolor f,g}\right]=0\iff-\textsc{kl}(q\mid\mid p)-\textsc{kl}(p\mid\mid q)=0\iff p=q.$ (26f)
@@
[…]
@@
We first establish conditions of $\alpha$ for which the variance of $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{cv}}}\right]$ does not increase that of $\mathop{\mathrm{Var}}\left[{\normalcolor\mu_{\textsc{mc}}}\right]$. In other words, we seek conditions on $\alpha$ such that the following inequality holds: